For a solution 0.10M of KOCN:

a. Write the hydrolisis reaction.
b. Write the equilibrium expression.
c. Find [H3O+], [OH-], pH, pOH, and [OCN-] for a solution 0-10 of potassium cianide.
Answer #1

a. KOCN (aq) + H2O (l)  \to HOCN (aq) + KOH (aq)

Or, OCN- (aq) + H2O (l)  \to HOCN (aq) + OH- (aq)

b. Kh =  \frac{[HOCN][OH^{-}]}{[OCN^{-}][H_2O]}

The concentration of H2O is very large and reagarded as constant.

Now, Kh = \frac{[HOCN][OH^{-}]}{[OCN^{-}]}

c.

Concentration KOCN HOCN OH
Initial C 0 0
Change -Cx +Cx +Cx
Equilibrium C - Cx Cx Cx

Kh = (Cx × Cx) /(C - Cx )

Now, Kh = Cx2/1-x

Or, 1 >> x , then, Kh = Cx2

or, x = √(Kh/C)

Or, Cx = √(Kh× C)

Now, Kh = \frac{Kw}{Ka}

From ICE table ; [OH-] = Cx

Or , [OH-]  = (\sqrt{}\frac{Kw \times C}{Ka})

Or, - log [OH-] = \frac{1}{2} ( - logKw + logKa + logC )

Or, pOH = \frac{1}{2} ( pKw - pKa - logC)

Or, pOH = 7 -  \frac{1}{2} (pKa + logC)

pH = 14 - {7 - \frac{1}{2} ( pKa + logC)}

Or, pH = 7 + \frac{1}{2} ( pKa + logC).

Given, C = 0.10 M

pKa of HOCN = 3.46

Then, pH = 7 + \frac{1}{2} ( 3.46 + log 0.1)

= 7 + \frac{1}{2} ( 3.46 -1)

= 8.23.

[H3O+] = 10-8.23 = 5.89 × 10-9 M.

Then,

pOH = 14 - 8.23 = 5.77.

Or, [OH-] = 10-5.77 = 1.7×10-6 M